If f(x) = |x-25| &
g(x) = f(f(x)) then for x > 50, g(x) is

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:Hello student, please find answer to your question
f(x) = |x-25|
g(x) = f(f(x))
g(x) = f(|x-25|)
g(x) = ||x-25|-25|
For x > 50
g(x) = |x-25-25|
g(x) = |x-50| = x - 50
g'(x) = 1